OpenStudy (anonymous):

find the second derivative f(x) = e^(-x^2/2) find the second derivative f(x) = e^(-x^2/2) @Mathematics

myininaya (myininaya):

\[f'(x)=(\frac{-x^2}{2})'e^{\frac{-x^2}{2}}\]

OpenStudy (anonymous):

\[e^{-x ^{2}/2}\]

OpenStudy (anonymous):

second derivative?

myininaya (myininaya):

find the derivative of the first derivative

myininaya (myininaya):

\[f''(x)=(\frac{-x^2}{2})''e^{\frac{-x^2}{2}}+(\frac{-x^2}{2})'(\frac{-x^2}{2})'e^{\frac{-x^2}{2}}\]

OpenStudy (anonymous):

alright now I have to set that equal to zero and solve for x

myininaya (myininaya):

why? you are trying to find second derivative are you looking for inflection points?

OpenStudy (anonymous):

Second derivative e^(-x^2/2) *(x^2 - 1)

OpenStudy (anonymous):

the interval on which the graph is concave up and concave down and inflection points